One-tailed test; variance is unknown, α = 0.05, n = 15


1
Expert's answer
2022-06-03T09:16:58-0400

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=14" degrees of freedom and the critical value for a left-tailed one-tailed test is "t_c = -1.76131."

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=14" degrees of freedom and the critical value for a right-tailed one-tailed test is "t_c = 1.76131."